# Upgrading Democracies with Fairer Voting Methods

### v1

@@ Abstract
Voting methods are instrumental design element of democracies. Citizens use them to express and aggregate their preferences to reach a collective decision. However, voting outcomes can be as sensitive to voting rules as they are to people’s voting choices. Despite the significance and inter-disciplinary scientific progress on voting methods, several democracies keep relying on outdated voting methods that do not fit modern, pluralistic societies well, while lacking social innovation. Here, we demonstrate how one can upgrade real-world democracies, namely by using alternative preferential voting methods such as cumulative voting and the method of equal shares designed for a proportional representation of voters’ preferences. By rigorously assessing a new participatory budgeting approach applied in the city of Aarau, Switzerland, we unravel the striking voting outcomes of fair voting methods: more winning projects with the same budget and broader geographic and preference representation of citizens by the elected projects, in particular for voters who used to be under-represented, while promoting novel project ideas. We provide profound causal evidence showing that citizens prefer proportional voting methods, which possess strong legitimacy without the need of very technical specialized explanations. We also reveal strong underlying democratic values exhibited by citizens who support fair voting methods such as altruism and compromise. These findings come with a global momentum to unleash a new and long-awaited participation blueprint of how to upgrade democracies.
@@ 1 Introduction
Reaching collective decisions that benefit all often remains an unfulfilled promise. While voting methods have been instrumental to adaptively govern societies, they are often based on outdated approaches to foster collective intelligence [1, 2, 3]. This may put democracies at risk instead of empowering them [4]. Over-relying on limited-choice majoritarian voting often results in promoting competition instead of cooperation, in marginalization of minorities instead of inclusion, in few winners and many losers [5, 4, 6, 7, 2, 3, 8]. Polarization, exacerbation of inequalities, and radicalization of the public sphere [9, 10] are prominent implications of voting methods that fail to foster fairness by design.
@@
The challenge of adopting democratic innovations. Proportional voting methods and collective decision making have attracted broad interest by inter-disciplinary research communities [11, 12, 13]. This includes the fields of social choice, computational social science, economics, complexity theory and political science. Nevertheless, adopting the developed methods in practice and changing the status quo is itself a long-standing and multi-faceted challenge [14]. It often comes as a ‘chicken and egg’ problem: difficulties in communicating alternative voting methods effectively and lack of evidence of how they function in real world cause risk aversion at the political level and limited will for change, which further maintains the lack of evidence, also in regards to the legitimacy perception of voters [15, 16, 17, 2, 10]. Such unprecedented complexity for democratic innovations requires rigorous scientific methods [17, 14, 10, 18]. A blueprint for shifting to fairer voting methods is urgently needed.
@@
Participatory budgeting as a blueprint for democratic innovations. In recent years, there has been quite a number of noteworthy “democratic innovations". These include citizen assemblies, government labs (“GovLabs"), open innovation (open access/data/source, creative commons), crowd sourcing/sensing/funding, hackathons, fablabs, makerspaces, and citizen science. While related to some of these recent trends, our article focuses on significant interventions in participatory budgeting. This is a process designed to distribute a public budget of a city council in a bottom-up way by engaging citizens in collaborative brainstorming sessions to propose their own projects, solutions, and ideas, on which they later vote [19]. Citizens may also be involved in the implementation of selected projects. Participatory budgeting offers a great potential for demonstrating grass-roots democratic innovations in real-world, which can be the blueprint to scale up further in other democratic processes. Started in Porto Alegre in 1989, participatory budgeting has since been adopted in many places across the globe, from remote communities in Kenya, Indonesia, and the Philippines to major cities such as Barcelona, Warsaw, Helsinki, Paris, and New York [20, 21]. Its uptake has been particularly significant in Europe, with a surge in participatory initiatives in recent years [19]. By 2019, Europe accounted for more than half of all participatory budgeting schemes worldwide, with more than 5,000 initiatives [21]. Despite this uptake, the impact of fairer voting methods on voting outcomes as well as project ideation and implementation are, so far, under-explored questions. This is despite their profound importance for the quality of the voting outcomes.
@@
Collective choice made fairer. We demonstrate a breakthrough for democracies by applying two alternative voting methods designed for a more expressive and proportional representation of citizens in a real-world setting. This has been achieved with a novel field test designed for a systematic assessment of a new participatory budgeting process in the city of Aarau in Switzerland. We used the ballot format of cumulative voting [22, 23] allowing voters to distribute 10 points over at least 3 out of a total of 33 project ideas. Compared to other prominent input voting methods such as approval, ranking, or knapsack voting [24], cumulative voting enables voters to express their preferences more effectively by weighting their support to different options. This expression of preferences is more inclusive and can provide a more proportional representation of minority groups. Complementary, we used the voting method of equal shares [25] to aggregate the points allocated by citizens and elect the winning project ideas. Equal shares promotes a more proportional representation in comparison with majoritarian voting. This is favorable, as majoritarian approaches to elect winners can sometimes lead to a "tyranny of the majority", where "the winner takes it all", i.e. monopolizes all the budget [26]. Proportional representation is achieved by assigning to every voter a decision power, which is given by an equal share of the budget. This manifests both in meeting voters’ preferences and in the geographical coverage of the winners [2]. In practice, the method of equal shares is resourceful, as it is likely to select a larger number of projects, which are cost-efficient, while satisfying the preferences of more people and increasing overall voter satisfaction [27, 28]. Given the applied principle of proportionality, we refer to these alternative decision-making approaches here as ‘fair voting methods’. As it turns out, they also promote legitimacy and inclusion. Our study provides universal findings, beyond Aarau, as we also study a large number of past participatory budgeting processes and recent follow ups adopting fair voting methods.
Overcoming barriers of shifting to fair voting methods. Combining cumulative voting and equal shares in a real-world participatory budgeting process allows us to unravel: (i) how to systematically adopt fair voting methods. Apparently, for policy-makers and a general public familiar with elections, the method of equal shares is pretty different from the widespread and longstanding standard method of utilitarian "greedy" voting, which simply sums up the votes and selects the most popular project ideas fitting the budget in a majoritarian way; (ii) how the method of equal shares influences participation in the project ideation phase, allowing for the consideration of additional criteria during the voting process. Project costs become instrumental, as more expensive projects require broader support. This may discourage expensive projects, which could be compensated by a larger number of more resource-efficient projects which are fairer distributed in space [27]; (iii) the legitimacy of the voting outcomes. Although the method of equal shares is designed to promote fairness, sacrificing popular expensive projects may be perceived as unfair by their supporters, who may be a considerable share of the population. Whether the selection of more alternative projects can compensate for the potential disappointment is key for the perceived legitimacy of fair voting methods; (iv) causal factors that promote a shift to fair voting methods. The crucial question is, whether there are any personal traits or societal values that support and accelerate democratic upgrades based on fair voting methods.
A systematic and rigorous approach to upgrade democracies with fair voting methods. We co-designed the participatory budgeting process of "Stadtidee" ("City Idea") in collaboration with the city of Aarau in Switzerland. It was conducted in 2023 with an ongoing project implementation. It is characterized by the following innovations (see Section 4.1 for more information): (i) the real-world application of the method of equal shares in combination with cumulative voting as the ballot format. Taken together, they enable a better, proportional representation of citizens’ preferences; (ii) an inclusive participation fostered throughout the entire process. The citizens proposed 161 ideas, out of which 33 were voted on by 1703 citizens, women in majority and children included, with 17 projects selected for implementation [29], see Table 1 for an outline; (iii) a survey study was conducted before and after voting. This allows us to rigorously assess the new process in depth over its entire lifetime. We are also able to assess the legitimacy of fair voting methods before and after voting and to profoundly understand the dynamics of voting outcomes that these methods generate, including causal and explanatory factors; (iv) a secure and privacy-preserving voting using and extending the Stanford open-source platform for participatory budgeting [30, 31]. This ensures transparency and trust when using fair voting methods. These four aforementioned innovations offer a strong foundation of validity to rigorously test fair voting methods. They provide a blueprint for testing upgrades of democracy, which have been missing so far.
@@ 2 Results
This articles presents three key results:
@@
1. Fair ballot aggregation results in voting outcomes with more winners, while achieving better representation of voters’ preferences and novel project ideas, also in space.
@@
2. Voters find the adopted ballot aggregation method of equal shares more legitimate, by preferring it over the majoritarian standard method and finding it fairer.
@@
3. The shift to a fairer voting method distinguishes supporting voters with strong democratic values from voters without these values. Democratic values that explain the preference for more proportional voting methods include altruism, opposition to costly popular projects, and support for more expressive and inclusive ballot formats.
2.1 Fair ballot aggregation has distinguishing multi-faceted qualities
Why the method of equal shares results in more winners. Figure 1a shows the 33 projects proposed for voting and the related voting results. The following key observations can be made: (i) Equal shares selects 17 winning projects, with more than half of the proposed projects winning. In contrast, the standard utilitarian greedy method would select 7 projects only, which would imply 26 losing projects, thereby creating a significant deficit in representing the broad range of citizen preferences. (ii) The project entitled ‘Wild bees’ paradise’ is an example of an expensive popular loser using the method of equal shares, which would win under the standard method. With 33.64% of voters selecting this project, it is the 3rd most popular project, which would consume 40% of the overall available budget. The method of equal shares selects several other, lower-cost projects instead, each of which may be less popular, but they address together the preferences of a broader population. (iii) The mean budget share of the proposed projects is 14.33%, which is 48.37% lower than a typical proposed project within 345 past participatory budgeting elections conducted with the standard method [32]. This illustrates the impact of ballot aggregation design already in the project ideation phase: it is more resourceful by encouraging lower-cost projects of different and often novel nature, with the potential for synergetic impact [27].
@@ 2.1 Fair ballot aggregation has distinguishing multi-faceted qualities
Figure 1: Compared to the standard utilitarian greedy method, which is majoritarian in nature, the number of winning projects resulting from the method of equal shares increases dramatically, with more than half of the proposed projects winning, while citizens’ proposed project ideas are better represented geographically. (A) Equal shares elects 17 winning projects compared to 7 with the majoritarian standard method. The project ‘Wild bees’ paradise’, requiring 40% of the total budget, is popular winner under the standard method but loser under equal shares. This allows more low-cost projects to win, which, in combination, better represent voters’ preferences. (B) Under equal shares, five districts receive more winning projects as compared to the standard method, thereby strengthening the periphery of the city. (C) Spatial fairness increases by 44% under equal shares as measured by the relative reduction in the index of dispersion for the number of winning projects relative to the proposed ones for each district.
@@
Equal shares improves voters’ geographic representation, especially for city peripheries. Figure 1b illustrates that the method of equal shares results in 5 districts (27.77%) with more winning projects compared to the majoritarian standard method, while no district receives a lower number of projects. This particularly benefits the periphery of the city that would otherwise be under-represented. Moreover, Figure 1c measures the relative number of winning/proposed projects in each district. Using the method of equal shares, the spatial fairness increases by 44% when measured by the reduction in the index of dispersion, which demonstrates that more proposed project win in each district. These results confirm that the periphery of Aarau is better represented, both in terms of the voter preferences and the number of proposed project ideas selected.
The method of equal shares represents voters better. Figure 2a shows the share of voters for different cumulative levels of the mean representation of their preferences. Representation is measured by the relative number of projects that are preferred by a voter and win. Comparing the standard method of 345 past participatory budgeting processes [33, 32] with the outcomes for the equal shares method in Aarau, we observe 3.7% more voters with higher representation levels, in particular for lower levels of cumulative representation. This means that equal shares better preserves the representation of the population even for under-represented voters, whose preferences do not align well with the voting outcomes. Specifically, for the participatory budgeting process in Aarau, the voters’ representation is 19% higher under equal shares compared to the standard method and 14.5% higher on average compared to 345 earlier participatory budgeting processes.

### v2

@@ Abstract
Voting methods are instrumental design elements of democracies. Citizens use them to express and aggregate their preferences to reach a collective decision. However, voting outcomes can be as sensitive to voting rules as they are to people’s voting choices. Despite significance and interdisciplinary scientific progress, several democracies keep relying on outdated voting methods that do not fit modern, pluralistic societies well, while lacking social innovation. Here, we demonstrate how one can upgrade real-world democracies, namely by using alternative preferential voting methods such as cumulative voting and the method of equal shares designed for a proportional representation of voters’ preferences. We rigorously evaluate the striking voting outcomes of these fair voting methods in a new participatory budgeting approach applied in the city of Aarau, Switzerland, including past and follow-up evidence. Results show more winning projects with the same budget. They also show broader geographic and preference representation of citizens by the elected projects, in particular for voters who used to be under-represented. We provide causal evidence showing that citizens prefer proportional voting methods, which possess strong legitimacy without the need of very specialized technical explanations. We also reveal strong underlying democratic values exhibited by citizens who support fair voting methods such as altruism and compromise. These findings come with the momentum to unleash a new and long-awaited participation blueprint of how to upgrade democracies globally.
@@ 1 Introduction
Reaching collective decisions that benefit the broader society often remains an unfulfilled promise. While voting methods have been instrumental to govern societies with citizens’ participation, they are often based on outdated approaches to foster collective intelligence [Kearns2009, Arnold2017, Mann2017]. This may put democracies at risk instead of empowering them [Emerson2020]. Standard majoritarian voting methods, for example, over-prioritize popular preferences over those of under-represented groups. Disregarding alternative voices and their value often results in lost opportunities for all. Over-relying on limited-choice majoritarian voting frequently results in promoting competition instead of cooperation, in marginalization of minorities instead of inclusion, in few winners and many losers [Santos2024, Emerson2020, Pournaras2020, Helbing2023, Arnold2017, Mann2017, Hanggli2021]. Polarization, exacerbation of inequalities, and radicalization of the public sphere [Lorenz2023, Bruter2017] are prominent implications of voting methods that fail to foster fairness by design. Therefore, the lack of fair, proportional representation of the different preferences in a society is a persistent setback to address.
@@
The challenge of adopting democratic innovations. Proportional voting methods and collective decision making have attracted a broad interest by interdisciplinary research communities [Pukelsheim2017, Lewis2018, Fisher2018]. This includes the fields of social choice, computational social science, economics, complexity theory, and political science. Nevertheless, adopting these methods in practice and changing the status quo is itself a long-standing and multi-faceted challenge [Madhavan2017]. It often comes as a ‘chicken and egg’ problem, (i) with difficulties in communicating alternative, fairer voting methods effectively and (ii) with a lack of evidence of how they function in the real world, which mutually enforce each other. This causes risk aversion at the political level and limited will for change, which further maintains the lack of evidence, also with regard to the legitimacy perception of voters [Cho2018, Park2023, Noveck2017, Arnold2017, Bruter2017]. Such unprecedented complexity for democratic innovations requires rigorous scientific investigation [Noveck2017, Madhavan2017, Bruter2017, Eliassi2020]. A blueprint for shifting to fairer voting methods is urgently needed.
@@
Participatory budgeting as a blueprint for democratic innovations. In recent years, there has been quite a number of noteworthy “democratic innovations". These include citizen assemblies, government labs (“GovLabs"), open innovation (open access/data/source, creative commons), crowd sourcing/sensing/funding, hackathons, fablabs, makerspaces, and citizen science. While related to some of these recent trends, our article focuses on novel interventions in participatory budgeting and their broader impact. This is a process designed to distribute some public budget of a city council in a bottom-up way by engaging citizens in collaborative brainstorming sessions to eventually propose their own projects, solutions, and ideas, on which they later vote [Fairstein2023]. Citizens may also be involved in the implementation of the selected projects.
How participatory budgeting generalizes. Participatory budgeting offers a great potential for demonstrating grassroots democratic innovations in the real world, which can be a blueprint that can be scaled up and wide later in other (collective) decision-making processes. For instance, the selection of projects to counter environmental problems, the distribution of research funds, the use of pensions funds, and entrepreneurial or stock market investments are all timely and pertinent budgeting processes that can benefit remarkably from budget-aware voting methods and, hence, from democratic innovations. Initiated in Porto Alegre back in 1989, participatory budgeting has since been adopted in many places across the globe, from remote communities in Kenya, Indonesia, and the Philippines to major cities such as Barcelona, Warsaw, Helsinki, Paris, and New York [Wampler2021, Dias2018]. Its uptake has been particularly significant in Europe, with a surge of participatory initiatives in recent years [Fairstein2023]. By 2019, Europe accounted for more than half of all participatory budgeting schemes worldwide, with more than 5,000 initiatives [Dias2018]. Despite the profound importance of voting methods for the quality of the outcomes [Arrow1964], the impact of fairer voting methods has, so far, been under-explored.
Collective choice made fairer. We demonstrate a breakthrough for democracies by applying alternative voting methods designed for a more expressive and proportional representation of citizens in real-world settings. This has been achieved with a novel approach for field-testing and systematic assessing a new participatory budgeting process in the city of Aarau, Switzerland, called “Stadtidee” (“City Idea”). The participatory budgeting process there was so successful that it is planned to repeat. It was also followed by similar implementations elsewhere (e.g. in Poland), which support our innovative approach and conclusions. The outcomes in Aarau are particularly stunning as compared to a similar “Stadtidee” format in Zurich, where the conventional majoritarian voting method was applied (see Section 2.2). Therefore, our field test is documented as a methodological breakthrough that is of real-world political relevance.
@@
We used the ballot format of “Cumulative Voting” [Cooper2007, Skowron2025] allowing voters to distribute 10 points over at least 3 out of 33 project ideas. Compared to other prominent input voting methods such as approval, ranking, or knapsack voting [Goel2019], cumulative voting enables voters to express their preferences more effectively by weighting their support for different options. This expression of preferences is more inclusive and can provide a more proportional representation of minority groups. Complementary, we used the “Method of Equal Shares” [Peters2021] to aggregate the points allocated by citizens and elect the winning project ideas. This voting method promotes a more proportional representation in comparison with majoritarian voting. Proportionality is the favorable aspect of fairness we study in this article. It avoids an undesirable feature of majoritarian voting approaches, which can sometimes lead to a "tyranny of the majority", where "the winner takes it all", i.e. monopolizes all the budget [Nyirkos2018]. In contrast, proportional representation is achieved by assigning to each voter the same decision power, which is given by an equal share of the budget to be distributed. This manifests both, in a better reflection of voter preferences and in a fairer geographical distribution of winning projects [Arnold2017]. In practice, the method of equal shares is resourceful, as it is likely to select a larger number of projects, which are cost-efficient, while satisfying the preferences of more people and increasing overall voter satisfaction [Maharjan2024, Yang2024] as shown in Section 2.1. Given the principle of proportionality that we applied, we refer to these alternative decision-making approaches here as “fair voting methods”. As it turns out, they also promote legitimacy and democratic values (see Section 2.3 and 2.4 respectively). Our study provides universal findings, beyond Aarau, as we also study a large number of past participatory budgeting processes and recent follow-ups adopting fair voting methods.
Overcoming barriers of shifting to fair voting methods. Combining cumulative voting (for casting the vote) and equal shares (for aggregating the points and electing the winning projects) in a real-world participatory budgeting process allows us to unravel:
  • (i)
    how to systematically adopt fair voting methods; apparently, for policy-makers and a general public familiar with elections, the method of equal shares is pretty different from the widespread and longstanding standard method of utilitarian “greedy” (i.e., majoritarian) voting, which simply sums up the votes and selects the most popular project ideas fitting the budget in a majoritarian way;
  • (ii)
    how the method of equal shares influences participation in the project ideation phase, allowing for the consideration of additional criteria during the voting process; importantly, project costs become instrumental, as more expensive projects require broader support, which may discourage expensive projects; this allows for a larger number of more resource-efficient projects which are fairer distributed in space [Maharjan2024];
  • (iii)
    the legitimacy of the voting outcomes; although the method of equal shares is designed to promote fairness, sacrificing popular expensive projects may be perceived as unfair by their supporters, who may be a considerable share of the population; whether the selection of more alternative projects can compensate for the potential dissatisfaction is key for the perceived legitimacy of fair voting methods;
  • (iv)
    causal factors that promote a shift to fair voting methods; the crucial question is, whether there are any personal traits or societal values that support and accelerate democratic upgrades based on fair voting methods.
A systematic and rigorous approach to upgrade democracies with fair voting methods. We co-designed the participatory budgeting process called “Stadtidee” (“City Idea”) in collaboration with the city of Aarau, Switzerland. The process was implemented in the year 2023. It was characterized by the following innovations (see Section 4.1 for more information):
  • (i)
    a real-world application of the method of equal shares in combination with cumulative voting as the ballot format; combined with each other, these methods enable a better, proportional representation of citizen preferences;
  • (ii)
    an inclusive participation fostered throughout the entire process; the citizens proposed 161 ideas, out of which 33 were voted on by 1703 citizens, women in majority and children included, with 17 projects selected for implementation [CityIdeaReport2025], see Table 1 for an outline;
  • (iii)
    a survey study was conducted before and after voting with a control condition; this allowed us to rigorously assess the new process in depth over its entire lifetime; we were also able to assess the legitimacy of fair voting methods before and after voting and to profoundly understand the dynamics of voting outcomes that these methods generate, including causal and explanatory factors;
  • (iv)
    a secure and privacy-preserving voting using and extending the Stanford open-source platform for participatory budgeting [SPBStadtidee2023, Wellings2023]; this ensures transparency and trust when using fair voting methods.
A comprehensive spatio-temporal evidence. To confirm and generalize the qualities of voting outcomes reached by fair voting methods, we highlight the continuum of evidence we provide:
  • (i)
    Local comparison: The actual use of equal shares in Aarau vs. the hypothetical use of the standard majoritarian method in Aarau (shown in Section 2.1);
  • (ii)
    Global comparison: The actual use of equal shares in Aarau vs. the actual use of the standard majoritarian method in past elections (shown in Section 2.2);
  • (iii)
    Confirmation on earlier evidence: The hypothetical use of equal shares in past elections vs. the actual use of the standard majoritarian method in past elections (shown in Section 2.2);
  • (iv)
    Confirmation on follow-up evidence: The actual use of equal shares in Aarau vs. the actual use of equal shares in follow-up elections of Aarau (shown in Section 2.2);
These four aforementioned innovations, in combination with comprehensive spatio-temporal comparisons with earlier and follow up real-world elections, enabled a rigorous test of fair voting methods and offers a strong foundation for the validity of our conclusions. This provides a promising blueprint for innovative upgrades of democracies in the future.
@@ 2 Results
This article presents three key results:
  • (1)
@@
    Fair ballot aggregation results in voting outcomes with more winners, while achieving better representation of voter preferences and diversity of project ideas, also in space (Section 2.1 and 2.2).
  • (2)
@@
    Voters find the adopted ballot aggregation method of equal shares more legitimate, preferring it over the majoritarian standard method and finding it fairer (Section 2.3).
  • (3)
@@
    The shift to a fairer voting method benefits from democratic values and promotes them. In fact, democratic values such as altruism, support for more expressive and inclusive ballot formats, and opposition to costly popular projects causally explain the preference for more proportional voting methods (Section 2.4).
2.1 Fair voting methods show multi-faceted qualities
Why the method of equal shares results in more winners. Figure 1a shows the 33 projects proposed for voting (very left column) and the related voting results (right columns under "Standard" and "Equal Shares"). The following key observations can be made:
  • (i)
    Equal shares selects 17 winning projects, with more than half of the proposed projects winning; in contrast, the standard majoritarian utilitarian greedy method would select 7 projects only, which would imply 26 losing projects, thereby creating a significant deficit in representing the broad range of citizen preferences;
  • (ii)
    The project entitled “Wild bees’ paradise” is an example of an expensive popular loser using the method of equal shares, which would win under the standard majoritarian method; with 33.64% of voters selecting this project, it is the 3rd most popular project, but it would consume 40% of the overall available budget. Similarly, the project“Open Children Studio" is the other expensive popular loser consuming 18.2% of the budget. The method of equal shares selects several other, lower-cost projects instead, each of which may be less popular, but together they better address the preferences of a broader population: 75% representation compared to 50% that “Wild bees’ paradise” and “Open Children Studio" achieve together.
  • (iii)
    The mean budget share of the proposed projects is 14.33%, which is 48.37% lower than a typical proposed project within 345 past participatory budgeting processes conducted with the standard majoritarian method [pabulib]; this illustrates the impact of ballot aggregation design already in the project ideation phase: it is more resourceful by encouraging lower-cost projects of different and often novel nature. This has the potential for synergetic impact creation due to the greater variety of cost-effective projects [Maharjan2024].
@@ 2.1 Fair voting methods show multi-faceted qualities
Figure 1: Compared to the utilitarian greedy ("Standard") method, which is majoritarian in nature, the number of winning projects resulting from the method of equal shares (see column "Equal Shares") increases dramatically, with more than half of the proposed projects winning, while the citizens’ proposed project ideas are better represented geographically. (A) Equal shares (s)elects 17 winning projects as compared to 7 with the majoritarian standard method. The project ‘Wild bees’ paradise’, requiring 40% of the total budget, is popular winner under the standard method, but a loser under equal shares. This allows more low-cost projects to win, which, in combination, represent the voter preferences better. (B) Under equal shares, five districts receive more winning projects as compared to the standard majoritarian method, thereby strengthening the periphery of the city. (C) Spatial fairness increases by 44% under equal shares. This is measured by the relative reduction in the index of dispersion for the number of winning projects relative to the proposed ones for each district.
@@
Equal shares improves the voters’ geographic representation, especially for city peripheries. Figure 1b illustrates that the method of equal shares results in 5 districts (27.77%) with more winning projects compared to the standard majoritarian method, while no district receives a lower number of projects. This particularly benefits the periphery of the city that would otherwise be under-represented. Moreover, Figure 1c measures the relative number of winning as compared to the number of proposed projects in each district. Using the method of equal shares, the spatial fairness increases by 44% when measured by the reduction in the index of dispersion, which demonstrates that more proposed project win in each district. These results confirm that the periphery of Aarau is better represented by the method of equal shares, both in terms of the voter preferences and the number of proposed project ideas selected.
2.2 Generalizing fair voting methods as a distinguishing paradigm
The method of equal shares represents voters better. Figure 2a shows the share of voters for different minimum levels of the mean representation of their preferences (i.e. what share of voters have at least 0.1, 0.2,…,0.9 representation). Representation is measured by the relative number of projects preferred (i.e. selected) by a voter among those that win. The mean is calculated first for the voters of each election and then among elections (Figure LABEL:fig:pabulib-representation-voters is generated by averaging all voters among all elections). This figure reveals whether we have few or many voters that are very satisfied or unsatisfied (high/low representation). To generalize our findings we compare the following:
  • (i)
    The standard method of 38/345 past participatory budgeting processes [Faliszewski2023, pabulib] (baseline) that share similar characteristics with the process in Aarau (matched, see Section LABEL:sec:choice-ballot-aggregation);
  • (i)
    The method of Equal shares applied hypothetically to these 38 past elections;
  • (i)
    The method of Equal shares as applied in Aarau.